Night #2 Soul Com - UT 30-Sep-2020

OSA: D.G. Huerta

ISp: G. Bechetti

Mtn Manager: B. Bingham

ISA: B. Rothberg (Secret Bunker)

AO: J. Power, G. Taylor (Tucson)


Plan:

  1. NGC 6240 - off axis
  2. OSCO - 1 or 2 sets of data
  3. HR 8799
  4. NGC 628 Region 1
  5. Performance Testing Bin 4
  6. Contingency: ESM for poor conditions

Problems:

  1. ADI seemed to work without any serious issues. There was more "slosh" in X of the bayside stages than in Y, but nothing too significant. The star did not wander off the edges
  2. DIMM is working in a hobbled mode. We get some data but it prefers super bright stars in the field. Thanks to S. Allanson for getting us as far as we did.

Night Log:

Prep:

UT 00;24 -starting up LUCI GUIs

UT 00:28 - taking test darks

L1 0001-0002 2.51 sec darks

UT 00:30 - J. Power starting up AO WFS

UT 01:17 - opening enclosure
NGC 6240:

UT 01:45 - looks like we are collimating OK...

Guider seeing estimate 0.7"

UT 01:46 - sending preset to NGC 6240

UT 01:47 - 0.8" guider, 0.7" WFS

UT 01:53 - will try gopt

Seeing on guider/wfs 0.69"//0.63"

OCAM mag 13.7, bin2, frequency 854, average pupil intensity 28, 5.5 photons/subapp (catalog Rmag = 12.15)

DIMM IS OPERATIONAL - WE LOVE YOU STEVE!

DIMM shows 0.89"

Active guiding on off-image - guider/wfs 0.65" (r-band filter)

Active guiding on off-image - 0.8" guider

AO Ref star checkflux = R =13.6

Active guiding off-image: guider 1.2"/WFS 0.86"

UT 02:23 - elevation = 40degrees

L1 0003-0023

UT 02:27 - sending preset to PSF star (pair separation to match off-axis of NGC 6240)

-script sent with AORef mag same as offaxis AORef in NGC 6240 script

Guider/WFS 1" before closing loops

Currently at Elevation 45 degrees

AO info: OCAM mag 14.3, bin2, frequency 858, average pupil intensity 15, 3.1 photons/subapp
AO Ref star from off-axis NGC 6240 is 90 degerees from galaxy (E), AO ref star relative to the PSF star is S (180 degrees) - need to rotate PSF star 90 degrees during reduction
L1 0024-0028

UT 02:42 - Completed

OSCO

UT 02:43 - getting ready for OSCO

Target is HD 230519

AO system is safe skipping AFTER the loops opened and AO stopped (leftover from previous target)

UT 03:04 - still working on preparing to run OSCO

UT 03:08 - starting OSCO

Pal 10

UT 03:15 - sending preset to Pal 10

loaded wrong version of script (no mask because we got 3 versions of this script with continual errors0
re-rotator error, need to change PA

UT 03:23 - sending new version with PA = 306 - had to choose a different guidestar

Guider WFS before closing loop: 0.55"/0.61"

AO info:

Catalog mag: 12.06, OCAM magnitude: 13.3, Bin 2, frequency 928, average pupil intensity 35, 7.1 photons/subapp
L1 0049 - looks like offset according within cube

UT 03:33 - Shell RIP'd - need to recover the adsec

L1 0049 - 0054 - do not use

UT 03:37 - sending preset again

AO info: Catalog mag: 12.06, OCAM magnitude: 13.4, Bin 2, frequency 914, average pupil intensity 33, 6.8 photons/subapp
Diagnosis of AO rip: AO wasn't pausing and resuming on offsets properly, tried to recocer while doing an offset, drove the forces too high which caused RIP'd

UT 04:05 - DIMM 0.92"

L1 0051-0077

UT 04:15 - Completed script
OSCO Part Deux

UT 04:16 - need to rest and reset the shell

UT 04:18 - shell is set

UT 04:34 - OSCO running

UT 04:39 - completed

HR 8799 - ADI Test

-Goal is to test the ADI issue near and at object transit at 79 elevation. Collection of science-quality data is NOT the primary goal

UT 04:41 - sending collimation script

UT 04:44 sending science script

-got telescope commit error on LUCI messages
-this was the Rmag error in the script - fixing now

838.1229

+2.79" DETX -3.075" DETY

-commiting offset

landed 1024,1024, Absorbing offset

subwin setup: 896,896 256 256

H2 image 1026,1020 - core is flat-topped

UT 05:00 - starting main part of script

-script stopped - needed to hit resume on AO

Do Not Have NCPA running

UT 05:06 - EL 74

UT 05:08 - sending DETXY absolute 0,0

L1 0107 - pixel position 1019.1017 (El 75)

AO info: Cataog mag: 5.94, OCAM magnitude: 7.3, Bin 1, frequency 1700, average pupil intensity 150, 7.3 photons/subapp

UT 05:16 - sending 0,0 absolute DETXY offset

L1 0114 - pixel position 1024, 1017 (El 76)

UT 05:23 sending 0,0 absolute DETXY offset

L1 0121 - pixel position 1019,1016 (El 77)

L1 0122 - also a 0,0 in script 1018,1016 (El 77)

UT 05:32 - sending 0,0 absolute DETXY offset

L1 0129 - pixel position 1025,1022 ( El 77)

UT 05;38 - loop paused - had to hit resume on AO

UT 05:40 sending 0,0 absolute DETXY offset

L1 0136 pixel position 1021, 1020 (El 78:20)

L1 0144 - 0,0 absolute offset from script pixel position 1019,1020 (El 78:32)

UT 05:59 - 0,0 absolute offset DETXY sent

L1 0156 - pixel position 1027,1020 (El 78:20 - other side of the transit)

L1 0165 - pixel position 1019,1022 (El 77:47)

-will go for 85 degrees of field rotation, another 15 minutes

UT 06:49 - sending 0,0 Absoluate DETXY offset

pixel position 1026,1032 (El 72:21)

Last frame L1 0207

L1 0208 - 128x128 subframe - flat topped

L1 0209 - 128x128 ND2 + FeII

L1 0210 - 256x256 ND2+FeII

L1 0211-0215 ND2+FeII same exposure time 256x256 (peaks 1800counts per DIT)

L1 0216-0220 ND2+H2 same exoposure time 256x256


NGC 628

UT 07:09 sending Region 2 script

UT 07:12 - kinda close to the moon - 40 degrees away

-taking some time to collmate

-optimize gain did not work do to high background from the galaxy

-LUCI timed out after trying to close the loops, attempted to continue at the insturment config and instead it resent the preset

-should go faster to acquire.

Seeing in guider is 0.5"/0.6"

L1 0220 - GEIR image looks odd and blasted

looks like either etched scrtatches or blank no counts, even when switching to LIR, restarted GEIRS aftter init all on filters and moving to Kband

- sending 0,0 absolute offset DETXY to get back to field,- looks like a script error they did a 15second MER x 4 saved in a cube, this seems to have gotten GEIRS in a funky state that even LIR wasn't working

UT 07:47 - -remade script but have to manually reconfigure for Br Gamma (left it in K)

-also had to align the FM4

UT 07:55 - restarting script

AO Info: Magnitude sent: 14.23, OCAM mag 15.8, bin 3, freq 588, average pupil intensity 15, 3.0 photons/subapp

First Exposure L1 0227-
-safe skip after a large offset

L1-0229 - shows a double image - DO NOT USE

inserted pause after L1 0229 to optimize gain

Restartng after the PSF pair of imagaes

guiding in off dither - 0.56"/0.58" in guider/WFS

UT 08:23 - guiding on off field 0.57" WFS

UT 08:34 - guiding on off field 0.62"/0.59" guider/WFS

UT 09:08 - guding on off field 0.72"/0.75" guider/WFS

L1 0257 - image quality bad - loops opened up - DO NOT USE

UT 09:35 - guiding on off exposure guider/WFS = 0.89"/0.63"

L1 0230-L0260

Completed at UT 09:38
Performance Testing

UT 09:42 - going to Star 5 02 59 45 +50 52 28 R=14.2

UT 09:45 - pointing check

UT 09:48 - sending preset to Star 5 again

Seeing 0.62"/0.57" guider/WFS

WFS reports Rmag=15.5

AO Info: sent magnitude 14.22, OCAM mag 15.5, bin 3, freq 770, 15 average pupil intensity, 2.8 photons/subapp

Re0optimized gain, info:
Optimize gain - no minimum reached for tip tilt and ho1 overestimated. Tweaking parameters for optimize gain

UT 10:02 - reoptimizing gain again (needed to update files and reload them)

UT 10:03 - changed parameters again, re-running optimize gain

-looks good, applying optical gain

LUCI timed out again

-test exposure 10sec ~ 21k counts

Faint companion can be seen

-moving +30" Detxy to get sky-frame

UT 10:26 - now testing NCPA

UT 10;36 - DIMM back up again

Eleavtion of DIMM 71, telescope is 69 elevation

DIMM went away

UT 10:42 DIMM back again

UT 10:45 - stopping AO to check guide camera

guider/WFS show 0.4"/0.5"

UT 10:49 - moving to Star #7 R=14.7 02:59:47 +51:07:52

AO Info: Target sent 14.61, OCAM reports 15.6 mag, bin 3 , freq 648, 16 average pupil intensity, 3.0 photons/subapp

UT 11:04 - sending offset +30" X in DETXY for background image
UT 11:09 - offsert back to field

UT 11:17 - running with NCPA on

UT 11:20 - DIMM shows 0.93"

-guider/WFS show 0.48"/0.5"

UT 11:24 - moving to Star #4 R=15.05

AO Info: Catalog mag 14.22, OCAM mag 15.5 , bin3, freq 774, 13 average pupil intensity, 2.7 photons/subapp

6700 counts in 2.51 sec

UT 11:33 - starintg with NCPA

--Oops sent to Star 5 again (B. Rothberg's error)

UT 11:36 - sending to the REAL Star #4 R=15.05

AO Info: Catalog mag 16.04, OCAM mag 16.4, bin 3 , freq 280, average pupil intensity 16, 3.4 photons/subapp

UT 12:10 - resending with AO Reg = 16.51 in script (manually changed from catalog value)
-resending the preset

AO Info:Same star sent as 16.51 to force bin4:, skipping check flux, OCAM 16.5 mag, 13.3 photons/subapp, Forced: bin4, freq 491
UT 12:20 - star PSF looks terrible. In -12 deg twilight right now.

UT 12:26 sending preset to AO843 R=10.4

UT 12:29 - resending preset again (some issue with the AO side)

err = Resmsg[3] = PresetAO {FLAO}[1282] - [AOException] presetAO: FATAL - Timeout waiting for techviewctrl.L.STATUS (-5001) TIMEOUT_ERROR
UT 12:32 - trying one more time....

UT 12:42 - subframed 256x256 on star 9k counts in peak

AO Info: sent mag 10.4, OCAM mag 11.5, bin1 freq 1075, 28 average pupil intensity, 6.4 photons/subapp

UT 12:49 - closing enclosure - no cals this morning to give fills and mask exchange time.

-- BarryRothberg - 29 Sep 2020
Topic revision: r3 - 01 Oct 2020, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback